# On the Adversarial Risk of Test Time Adaptation: An Investigation into Realistic Test-Time Data Poisoning

### v2

@@ A.1.3 Detailed Poisoning & Training Algorithm
            $\mathcal{L}=loss_{1}+\sum^{L}_{l}\lambda_{l}\cdot loss_{2}^{l}$
@@
Algorithm 1 The pipeline of the proposed test time data poisoning
A.1.4 Details of Defense Practices
Here, we provide the details about the defense practices evaluated in Tab. 5 of the manuscript.
A.2 Additional Details for Experiment
Benchmark Poisoning Methods: We evaluated the following methods for test-time data poisoning. DIA (Wu et al., 2023) is one of the first approaches towards test-time data poisoning, generating the poisoned data via maximizing the cross-entropy of other benign data. We adapt DIA to the realistic evaluation protocol by splitting $\mathcal{B}_{ab}$ into two subsets of $50\%$ each, i.e. $\mathcal{B}^{p}_{ab}:\mathcal{B}^{b}_{ab}=1:1$, and craft $\mathcal{B}^{p}_{a}$ to maximize the cross-entropy of $\mathcal{B}^{b}_{ab}$. TePA (Cong et al., 2023) proposed to maximize entropy to generate poisoned data and performed attack before TTA starts. We adapt TePA to generate the poisoned data based on the source model and inject them into TTA pipeline on-the-fly. MaxCE (Goodfellow et al., 2014) is an established way to create adversarial samples by maximizing the cross-entropy loss. Finally, we evaluate the two attack objectives proposed in this paper, i.e high entropy attack (NHE Attack) and low entropy attack against most probable and balanced non ground-truth class (BLE Attack). For both NHE Attack and BLE Attack, we evaluate the attack objective subject to the constraint of our proposed feature consistency (Eq. 6).
@@ A.2 Additional Details for Experiment
Datasets: We evaluate on three datasets, widely adopted for TTA benchmarking. CIFAR10-C, CIFAR100-C and ImageNet-C are synthesized from the original clean validation set by adding various types of corruptions to simulate natural distribution shifts (Hendrycks & Dietterich, 2019). We choose corruption level 5 and perform continual test-time adaptation setting for evaluation. Following prior works (Wang et al., 2022; Döbler et al., 2023), we adopt the pre-trained WideResNet-28 Zagoruyko & Komodakis (2016), ResNeXt-29 (Xie et al., 2017), and ResNet-50 (He et al., 2016) models for experiments on the CIFAR10-C, CIFAR100-C, and ImageNet-C datasets, respectively.
@@
A.3 Additional Empirical Analysis
A.3.1 Visualization of Poisoned Samples
A.3.2 Different Attack Budgets
@@
We further evaluate the effectiveness of proposed poisoning approach under different attack budgets. Specifically, we evaluated at $b=0.1$, $b=0.2$ and $b=0.5$. We clearly observe that both high entropy and low entropy attacks are effective regardless of attack budgets.

### v3

@@ A.1.6 Detailed Poisoning & Training Algorithm
            $\mathcal{L}=loss_{1}+\frac{1}{L}\sum^{L}_{l}\lambda_{l}\cdot loss_{2}^{l}$
@@
Algorithm 1 The pipeline of RTTDP
A.2 experimental setups of DIA, TePA and RTTDP
We revisit the experimental setups of the previous methods, i.e. DIA (Wu et al., 2023) and TePA Cong et al. (2023), explain the differences under our RTTDP protocol, and justify the adaptations we made to ensure fair comparisons.
Commonalities among the different protocols: The three protocols, TePA, DIA, and RTTDP, share several overarching goals and assumptions. First, all protocols aim to evaluate the adversarial risks posed to Test-Time Adaptation (TTA) by injecting poisoned samples into the test data stream. Second, all protocols allow the adversary to obtain the source model, since the source model is usually the well-known pre-trained model, e.g., ImageNet pre-trained ResNet, and the open-source foundation model, e.g., DINOv2, SAM.
Key Differences among Protocols: Despite sharing some commonalities, the protocols diverge significantly in their attack setups.
In the TePA protocol, poisoned samples are generated by maximizing the entropy of the adversary’s crafted samples with respect to the source model’s predictions. These poisoned samples are injected into the TTA pipeline before any benign users’ samples are processed, simulating an offline attack scenario. However, this approach is unrealistic in real-world settings, where adversaries cannot fully control the sequence of test samples in advance.
In contrast, the DIA protocol generates poisoned samples by optimizing them to maximize the cross-entropy loss of benign samples belonging to other users. DIA assumes direct access to the online model’s parameters and the ability to observe other users’ benign samples. Poisoned samples are injected into the TTA pipeline alongside the corresponding benign users’ samples. However, this protocol has significant limitations in realistic settings. In practice, adversaries typically lack access to or control over the online model’s parameters. Additionally, it is highly improbable for adversaries to observe benign users’ samples, let alone the validation samples required for optimizing poisoning objectives.
The proposed RTTDP protocol addresses these limitations by operating under more realistic assumptions. In RTTDP, the adversary neither has access to other users’ benign samples nor the parameters of the online model. Instead, RTTDP employs a surrogate model, initialized as the source model, to generate poisoned samples. This surrogate model is iteratively updated based on feedback from previously injected poisoned samples. Poisoned samples are then injected into the TTA pipeline, either uniformly or non-uniformly, depending on the attack frequency in RTTDP protocol.
Adaptations of Competing Methods to RTTDP protocol: To ensure fair comparisons under RTTDP protocol, we made the following adjustments to the competing methods:
For TePA method, we preserved TePA’s original poisoning objective, i.e. maximizing entropy, but adapted the poisoned data injection strategy from an offline manner to an online manner, i.e. placing poisoning in between benign testing batches according to RTTDP.
For DIA method, (1) Replacing Online Model Parameters: DIA’s original objective relies on online model parameters, which are inaccessible in RTTDP. We replaced these parameters with the initial surrogate model, i.e. source model. (2) No Access to Benign Users’ Samples for Optimization: DIA uses benign users’ samples as optimization targets in its original setup. To meet RTTDP’s constraints, we split $\mathcal{B}_{ab}$ into two equal subsets, $\mathcal{B}_{ab}^{p}$ and $\mathcal{B}_{ab}^{b}$, with a 1:1 ratio. We then generate poisoned samples $\mathcal{B}_{a}^{p}$ by maximizing the cross-entropy loss of $\mathcal{B}_{ab}^{b}$. The specific formula can be found in Eq. 15.
A.3 Additional Details for Experiment
@@
Implementation Details: For a fair comparison, we implement various data poisoning methods within a unified poisoning framework. This framework utilizes a 40-step Projected Gradient Descent (PGD Boyd & Vandenberghe (2004)) optimization process tailored to the respective objectives of each method. The poisoned samples generated are then injected into the TTA pipeline in an online manner, adhering to the RTTDP protocol. Specifically, the respective objectives of different competing poisoning methods are shown as follows,
  • DIA (Wu et al., 2023):
    $\mathcal{B}^{p}_{a}=\arg\min_{\mathcal{B}_{a}^{p}}E_{(x,y)\in\mathcal{B}_{ab}^{b}}\left[-CrossEntropyLoss(h(x;\hat{\theta}_{0}(\mathcal{B}_{ab}^{b}\cup\mathcal{B}_{a}^{p})),y)\right]$ (15)
  • TePA (Cong et al., 2023):
    $\mathcal{B}_{a}=\arg\min_{\mathcal{B}_{a}}E_{(x,y)\in\mathcal{B}_{a}}\left[-Entropy(h(x;\hat{\theta}_{0}(\mathcal{B}_{a})))\right]$ (16)
  • MaxCE (Madry et al., 2018):
    $\mathcal{B}_{a}=\arg\min_{\mathcal{B}_{a}}E_{(x,y)\in\mathcal{B}_{a}}\left[-CrossEntropyLoss(h(x;\hat{\theta}_{0}(\mathcal{B}_{a})),y)\right]$ (17)
  • Unlearnable Examples (Huang et al., 2021):
    $\mathcal{B}_{a}=\arg\min_{\mathcal{B}_{a}}E_{(x,y)\in\mathcal{B}_{a}}\left[CrossEntropyLoss(h(x;\theta_{init}(\mathcal{B}_{a})),y)\right]$ (18)
  • Adversarial Poisoning (Fowl et al., 2021):
    $\mathcal{B}_{a}=\arg\min_{\mathcal{B}_{a}}E_{(x,y)\in\mathcal{B}_{a}}\left[CrossEntropyLoss(h;\hat{\theta}_{0}(\mathcal{B}_{a}),\hat{y})\right],\text{where}\ \ \hat{y}=(y+1)\%K.$ (19)
  • NHE Attack (Ours):
    $\mathcal{B}_{a}=\arg\min_{\mathcal{B}_{a}}\max_{\lambda}E_{(x,y)\in\mathcal{B}_{a}}\left[\mathcal{L}^{NHE}_{atk}(x;\hat{\theta}_{t}(\mathcal{B}_{a}))+\lambda\mathcal{L}_{reg}\right]$ (20)
  • BLE Attack (Ours):
    $\mathcal{B}_{a}=\arg\min_{\mathcal{B}_{a}}\max_{\lambda}E_{(x,y)\in\mathcal{B}_{a}}\left[\mathcal{L}^{BLE}_{atk}(x;\hat{\theta}_{t}(\mathcal{B}_{a}))+\lambda\mathcal{L}_{reg}\right]$ (21)
where $\hat{\theta}_{0}$ indicates the initial surrogate model parameters, $\theta_{init}$ indicates the randomly initialized parameters and $K$ is the number of the category in the dataset. Since under the RTTDP protocol the real-time model parameters are unavailable to access, we leverage the initial surrogate model, to replace the online model they might have used in their original paper, as the threat model for the competing methods. The surrogate model is initialized as source model. For our proposed methods, we employ the proposed surrogate model distillation module to update the surrogate model, and during each PGD iteration, the variables $\mathcal{B}_{a}$ and $\lambda$ are updated simultaneously using gradient descent for $\mathcal{B}_{a}$ and gradient ascent for $\lambda$, respectively. More details about our proposed methods can be found in Alg. 1.
A.4 Additional Empirical Analysis
A.4.1 Comparison with Adversarial Attack Methods
Adversarial attack methods are designed to generate perturbations on input samples to mislead the model into making incorrect predictions. Here, we aim to investigate whether the adversarial effects of poisoned samples, generated using advanced adversarial attack methods (Madry et al., 2018; Croce & Hein, 2020; Chen et al., 2022), can be effectively transferred to benign users’ samples under the RTTDP protocol.
We conduct the experiments on CIFAR10-C and ImageNet-C datasets with a Uniform attack frequency under our proposed RTTDP protocol. The results are shown in Tab. 8 and Tab. 9. We make the following observations. (i) The objectives of adversarial attack methods and data poisoning methods differ fundamentally. Adversarial attack methods focus on generating adversarial noise to mislead model predictions on the perturbed test samples. In contrast, data poisoning methods aim to inject carefully crafted poisoned samples to degrade the model’s performance on subsequent benign samples after adaptation. (ii) While AutoAttack (Croce & Hein, 2020) represents a more advanced adversarial attack method, its performance is inferior to that of MaxCE-PGD (Madry et al., 2018) on ImageNet-C. (iii) Furthermore, certain complex adversarial attack methods, such as GMSA-MIN and GMSA-AVG (Chen et al., 2022), require generating adversarial perturbations separately for each class, a process that incurs substantial computational costs and limits scalability (which is why these methods are excluded from comparison on ImageNet-C), and still fall short compared to the efficacy of our proposed data poisoning methods.
Table 8: Comparison with different adversarial attack methods with our proposed data poisoning methods on CIFAR10-C dataset under the RTTDP protocol.
Attack Objective TENT EATA SAR ROID Avg
NoAttack 19.72 18.03 18.94 16.37 18.27
MaxCE-PGD (Madry et al., 2018) 18.55 18.17 19.50 18.57 18.70
AutoAttack (Croce & Hein, 2020) 26.29 19.12 19.56 18.67 20.91
GMSA-MIN (Chen et al., 2022) 35.92 22.78 19.99 18.65 24.33
GMSA-AVG (Chen et al., 2022) 38.80 21.89 19.95 18.51 24.79
BLE Attack (Ours) 54.07 45.20 26.80 19.06 36.28
NHE Attack (Ours) 73.86 29.73 24.56 17.00 36.29
Table 9: Comparison with different adversarial attack methods with our proposed data poisoning methods on ImageNet-C dataset under the RTTDP protocol.
Attack Objective TENT SAR CoTTA ROID Avg
NoAttack 63.49 61.26 63.02 53.42 60.30
MaxCE-PGD (Madry et al., 2018) 62.64 61.66 68.83 59.89 63.26
AutoAttack (Croce & Hein, 2020) 64.48 61.42 63.03 54.78 60.93
BLE Attack (Ours) 68.04 64.31 66.40 57.10 63.96
NHE Attack (Ours) 78.03 72.58 66.40 57.72 68.68
A.4.2 Ablation study on query counts
Regarding varying query attempts, we add an additional evaluation as follows. Nonetheless, we want to highlight that the query attempts do not have to be limited for our method because all queries are submitted to the surrogate model rather than the online model. More queries simply makes generating poisoning slower. In this study, we vary the query steps from 10 to 60 for the projected gradient descent optimization (Boyd & Vandenberghe, 2004). We evaluate varying attack query counts for two TTA methods under their respective strongest attack objectives. The results in the Tab. 10 are obtained on CIFAR10-C dataset with a Uniform attack frequency. We make the following observations. (i) Increasing the number of queries could improve the performance at a low query budget. (ii) When the budget is increased to beyond 40 queries, the performance saturates. We draw the conclusion that allowing sufficient queries to the surrogate model is necessary for generating effective data poisoning, and, importantly, this procedure will not create alert to the online model.
Table 10: The ablation study on query counts. These results are obtained on CIFAR10-C dataset with a Uniform attack frequency under RTTDP protocol. We choose 40 queries throughout the experiments.
TTA Method 10 20 30 40 50 60
TENT (NHE Attack) 66.95 74.36 74.34 73.86 73.51 73.66
EATA (BLE Attack) 35.73 39.70 42.36 45.20 45.99 45.89
A.4.3 Analysis on Symmetric KLD used for distilling surrogate model
In this work, we adopt the common practice of symmetrizing the Kullback-Leibler Divergence (KLD) to ensure balanced alignment between distributions in the surrogate model distillation. Following the definitions provided in the main text, the forward KLD is expressed as $KLD(h(x_{i};\theta_{t})\|h(x_{i};\hat{\theta}_{t}))$, while the reverse KLD is defined as $KLD(h(x_{i};\hat{\theta}_{t})\|h(x_{i};\theta_{t}))$.
Forward KLD emphasizes penalizing discrepancies where the distilled (surrogate) model $\hat{\theta}_{t}$ assigns low probability to samples that the source (real-time target) model $\theta$ deems important. It encourages the distilled model to mimic the behavior of the target model by focusing on areas of high confidence in $\theta$’s posterior.
Reverse KLD, in contrast, focuses on matching $\theta$’s predictions where $\hat{\theta}$ assigns high probabilities. This can result in sharper, more focused distributions but might dismiss less probable regions of $\theta$’s posterior.
The symmetric KLD balances the above two objectives. The forward KLD may be more suitable when surrogate model is significantly smaller than the target model and the objective is to allow the surrogate model to mimic the target model’s certainty. When the surrogate model is of the same capacity with target model, using the symmetric KLD may better align the two models in both high confident and low confident predictions. In this work, the capacity of surrogate is similar to target model. Thus, we hypothesize that the symmetric KLD could be better.
We further use empirical observations in the Tab. 11 below to support the hypothesis. With symmetric KLD the performance is slightly better than using the forward KLD.
Table 11: Comparison between Symmetric KLD and Forward KLD used for surrogate model distillation. These results are obtained on CIFAR10-C dataset with a Uniform attack frequency under RTTDP protocol.
TTA Method Symmetric KLD $KLD(h(x_{i};\theta_{t})||h(x_{i};\hat{\theta}_{t}))$
TENT (NHE Attack) 73.86 74.35
EATA (BLE Attack) 45.20 43.99
SAR (BLE Attack) 26.80 26.35
Nevertheless, we do acknowledge that both symmetric KLD and forward KLD give competitive results. The choice depends on computation affordability and empirical observations.
A.4.4 Different Attack Budgets
@@
We further evaluate the effectiveness of proposed poisoning approach under different attack budgets. Specifically, we evaluated at $r=0.1$, $r=0.2$ and $r=0.5$. We clearly observe that both high entropy and low entropy attacks are effective regardless of attack budgets.